OpenStudy (anonymous):

3. Two ocean beaches are being affected by erosion. The table shows the width, in feet, of each beach at high tide measured where 1995 is represented by year 0. A. Describe the patterns shown by the erosion data measurements shown for each of the beaches in the table. B. Between which years will the beaches have approximately the same width? C. Assuming these rates remain constant, what can you do to get a better approximation of when the two beaches will have the same width?

OpenStudy (anonymous):

@Jonathan1997

OpenStudy (anonymous):

2.5x – 10.5 = 64(0.5x) 2.5x - 10.5 = 32x 2.5x - 32x = 10.5 -29.5x = 10.5 x = -10.5/29.5

OpenStudy (anonymous):

@Jonathan1997

OpenStudy (anonymous):

do you understand how to complete the questions?

OpenStudy (anonymous):

hi isaac